# Fast Gradient Computation for RoPE Attention in Almost Linear Time

### v2

@@ 4 RoPE Attention Gradient Calculation
  • The column function $u(x)_{j_{0}}\in\mathbb{R}^{n}$ (Definitions 3.12),
  • $\alpha(x)_{j_{0}}$ is a real number (Def. 3.13),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. 3.14),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. 3.16), and
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. 3.17).
@@
and we note that the 1st equality is by Def. 3.12, the 2nd equality is by chain rule, the 3rd equality is due to Part 1, and the 4th equality is because of Def. 3.12.
@@
and we note that the 1st equality is by Def. 3.13, the 2nd equality is by product rule, the 3rd equality is due to $\frac{\mathrm{d}\mathbf{1}_{n}}{\mathrm{d}x_{i}}={\bf 0}_{n}$, the 4th equality is because of Def. 3.12, and 5th equality derives from basic algebra.
@@
and we note that the 1st equality is by Def. 3.14, the 2nd equality is by product rule, the 3rd equality is due to chain rule, the 4th equality is because of previous parts, the 5th and 6th equalities derive from Def. 3.14.
@@
and we note that the 1st equality is by Def. 3.16, the 2nd equality is by $\frac{\mathrm{d}E_{j_{0},i_{0}}}{\mathrm{d}x_{i}}={\bf 0}_{n}$, and the 3rd equality is due to the previous part.
@@
and we note that the 1st equality is by Def. 3.17, the 2nd equality is by chain rule and the 3rd equality is due to the previous part. ∎
5 Exact Gradient Computation Time
In this section, we analyze the time complexity of exact gradient computation. In Section 5.1, we reformulate the closed form of the gradient. In Section 5.2, we show the time complexity for $s(x)$ and $v(y)$. In Section 5.3, we show the time complexity for $\ell(x)$. In Section 5.4, we show the time complexity for $\beta(x)$ and $\gamma(x)$. In Section 5.5, we show the total time complexity for computing the gradient of RoPE attention.
5.1 Reformulate the Gradient into Its Closed Form
@@
Lemma 5.1 (Gradient Reformulation, $\frac{\mathrm{d}\mathsf{Loss}(x)_{j_{0},i_{0}}}{\mathrm{d}x}$).
@@
  • The column function $u(x)_{j_{0}}\in\mathbb{R}^{n}$ (Definitions 3.12),
  • $\alpha(x)_{j_{0}}$ is a real number (Def. 3.13),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. 3.14),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. 3.16), and
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. 3.17).
@@
where the 1st step follows from Lemma 4.1, and all other steps follow from Fact 3.9.
@@
Lemma 5.2 (Gradient Reformulation, $\frac{\mathrm{d}\mathsf{Loss}(x)_{j_{0}}}{\mathrm{d}x}$).
@@
  • The column function $u(x)_{j_{0}}\in\mathbb{R}^{n}$ (Definitions 3.12),
  • $\alpha(x)_{j_{0}}$ is a real number (Def. 3.13),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. 3.14),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. 3.16),
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. 3.17), and
@@
$\frac{\mathrm{d}\mathsf{Loss}(x)_{j_{0}}}{\mathrm{d}x}=\widetilde{\mathsf{A}}_{j_{0}}^{\top}(s(x)_{j_{0}}\circ A_{3}\beta(x)_{j_{0}})-\widetilde{\mathsf{A}}_{j_{0}}^{\top}s(x)_{j_{0}}\langle s(x)_{j_{0}},A_{3}\beta(x)_{j_{0}}\rangle.$
@@
$\frac{\mathrm{d}\mathsf{Loss}(x)_{j_{0}}}{\mathrm{d}x}=$ $~{}\sum_{i_{0}\in[d]}\frac{\mathrm{d}\mathsf{Loss}(x)_{j_{0},i_{0}}}{\mathrm{d}x}$
@@
Lemma 5.3 (Gradient Reformulation, $\frac{\mathrm{d}\mathsf{Loss}(x)}{\mathrm{d}x}$).
@@
  • The column function $u(x)_{j_{0}}\in\mathbb{R}^{n}$ (Definitions 3.12),
  • $\alpha(x)_{j_{0}}$ is a real number (Def. 3.13),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. 3.14),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. 3.16),
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. 3.17),
@@
5.2 Time Complexity for Computing $s(x)$ and $v(y)$ Functions
In this section, we use the vector and matrix multiplication time complexity from Definition 3.5 and Fact 3.6 to analyze the complexity of computing $s(x)$ and $v(y)$.
Lemma 5.4.
Pick $s(x)$ and $v(y)$ from Def. 3.14 and Def. 3.15, then it costs $O(\mathcal{T}_{\mathrm{mat}}(n,d,d)+\mathcal{T}_{\mathrm{mat}}(n,d,n))$ time to get $s(x)$, and it costs $\mathcal{T}_{\mathrm{mat}}(n,d,d)$ time to get $v(y)$.
@@ 5.2 Time Complexity for Computing $s(x)$ and $v(y)$ Functions
5.3 Time Complexity for Computing $\ell(x)$ Functions
In this section, we use the vector and matrix multiplication time complexity from Definition 3.5 and Fact 3.6 to analyze the complexity of computing $\ell(x)$.
Lemma 5.5.
We have $\ell(x)$ from Def. 3.16, then it costs $\mathcal{T}_{\mathrm{mat}}(n,n,d)+O(nd)$ to calculate $\ell(x)$.
@@ 5.3 Time Complexity for Computing $\ell(x)$ Functions
Then, it requires $O(nd)$ time to get $s(x)v(y)-E$.
@@
5.4 Time Complexity for Computing $\beta(x)$ and $\gamma(x)$ Functions
In this section, we use the vector and matrix multiplication time complexity from Definition 3.5 and Fact 3.6 to analyze the complexity of computing $\beta(x)$ and $\gamma(x)$.
Lemma 5.6.
@@ 5.4 Time Complexity for Computing $\beta(x)$ and $\gamma(x)$ Functions
5.5 Time Complexity for Computing the Gradient of RoPE Attention
Finally, We are able to present the time complexity for computing the gradient of RoPE attention.
Theorem 5.7 (RoPE attention gradient computation time complexity).
@@ 5.5 Time Complexity for Computing the Gradient of RoPE Attention
    We need time $O(\mathcal{T}_{\mathrm{mat}}(n,d,d)+\mathcal{T}_{\mathrm{mat}}(n,d,n))$ for $s(x),v(y)$ from Lemma 5.4.
@@
    We need time $O(\mathcal{T}_{\mathrm{mat}}(n,n,d)+\mathcal{T}_{\mathrm{mat}}(n,d,d))$ for $\ell(x)$ from Lemma 5.5.
@@
    We need time $O(\mathcal{T}_{\mathrm{mat}}(n,n,d))$ for $\beta(x)$ from Lemma 5.6.
@@
    We need time $O(n^{2})$ for $\gamma(x)$ from Lemma 5.6.
@@
6 Low Rank Approximation of Attention
This section presents the fast running time using the low-rank approximations where the low-rank matrices are generated from the polynomial method (see Lemma 3.4).
Lemma 6.1 (Theorem 1.3 from [7]).
6.1 Approximate $s$ Using Low Rank Approximation
@@
Lemma 6.2 (Low Rank Approximate $s(x)$).
@@
6.2 Approximate $\ell$ Using Low Rank Approximation
@@
Lemma 6.3 (Low Rank Approximate $\ell(x)$).
Let $d$ equal $O(\log n)$. Suppose we can use $O(\log n)$ bits to write every entry in $E,v(y)\in\mathbb{R}^{n\times d}$. Define the $\ell(x)\in\mathbb{R}^{n\times d}$ as specified in Def. 3.16. Then, we have $U_{1},V_{1}\in\mathbb{R}^{n\times k_{1}}$ such that $\|U_{1}V_{1}^{\top}v(y)-E-\ell(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n).$
@@
where the 1st is because of Def. 3.16, 2nd step is based on the distributive law, and 3rd step is due to Lemma 6.2. ∎
6.3 Approximate $\beta$ Using Low Rank Approximation
@@
Lemma 6.4 (Low Rank Approximate $\beta(x)$).
Let $k_{2}=n^{o(1)}$. We define $\ell(x)\in\mathbb{R}^{n\times d}$ based on Def. 3.16, and $v(y)\in\mathbb{R}^{n\times d}$ based on Def. 3.15. We suppose $\beta(x)$ is equal to $v(y)\ell(x)^{\top}$, which is an $n\times n$ matrix. Let $U_{2},V_{2}\in\mathbb{R}^{n\times k_{2}}$ such that $\|U_{2}V_{2}^{\top}-\beta(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n)$. In $n^{1+o(1)}$ time, we can get $U_{2},V_{2}$.
@@
By Lemma 6.3, $U_{1}V_{1}^{\top}v(y)-E$ approximately equals to $\ell(x)$.
@@
where the first step is based on the definition of $\beta(x)$ and $\widetilde{\beta}(x)$, the second step is due to the distributive law, and the third step derives from Lemma 6.3.
@@
6.4 Approximate $\gamma$ Using Low Rank Approximation
@@
Lemma 6.5 (Low Rank Approximate $\gamma_{1}(x)$).
@@
Then, we have the following by Fact 3.8.
@@
where the first inequality is because of the def. of $\gamma_{1}(x)$, the second equality is due to the def. of $U_{3},V_{3}$, the third equality is based on Fact 3.8, the fourth equality is due to the def. of $\widetilde{s}(x)$ and $\widetilde{\beta}(x)$, the fifth equality is due to simple arithmetic, the sixth inequality is because of the triangle inequality, and the seventh inequality derives from Lemma 6.2 and Lemma 6.4. ∎
@@
Lemma 6.6 (Low Rank Approximate $\gamma_{2}(x)$).
@@
Then, we define $\widetilde{\rho}(x)_{j_{0}}$ as the inner product of $\widetilde{s}(x)_{j_{0}}$ and $\widetilde{\beta}(x)_{j_{0}}$, and by Fact 3.9, we have $\widetilde{\rho}(x)_{j_{0}}=(U_{1}V_{1})_{j_{0},*}\cdot(U_{2}V_{2})_{j_{0},*}^{\top}$
@@
6.5 Fast Computation in Almost Linear Time

### v3

@@ Appendix C RoPE Attention Gradient Calculation
  • The column function $u(x)_{j_{0}}\in\mathbb{R}^{n}$ (Definitions B.1),
  • $\alpha(x)_{j_{0}}$ is a real number (Def. B.2),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. B.3),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. B.5), and
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. B.6).
@@
and we note that the 1st equality is by Def. B.1, the 2nd equality is by chain rule, the 3rd equality is due to Part 1, and the 4th equality is because of Def. B.1.
@@
and we note that the 1st equality is by Def. B.2, the 2nd equality is by product rule, the 3rd equality is due to $\frac{\mathrm{d}\mathbf{1}_{n}}{\mathrm{d}x_{i}}={\bf 0}_{n}$, the 4th equality is because of Def. B.1, and 5th equality derives from basic algebra.
@@
and we note that the 1st equality is by Def. B.3, the 2nd equality is by product rule, the 3rd equality is due to chain rule, the 4th equality is because of previous parts, the 5th and 6th equalities derive from Def. B.3.
@@
and we note that the 1st equality is by Def. B.5, the 2nd equality is by $\frac{\mathrm{d}E_{j_{0},i_{0}}}{\mathrm{d}x_{i}}={\bf 0}_{n}$, and the 3rd equality is due to the previous part.
@@
and we note that the 1st equality is by Def. B.6, the 2nd equality is by chain rule and the 3rd equality is due to the previous part. ∎
C.1 Reformulate the Gradient into Its Closed Form
@@
Lemma C.2 (Gradient Reformulation, $\frac{\mathrm{d}\mathsf{Loss}(x)_{j_{0},i_{0}}}{\mathrm{d}x}$).
@@
  • The column function $u(x)_{j_{0}}\in\mathbb{R}^{n}$ (Definitions B.1),
  • $\alpha(x)_{j_{0}}$ is a real number (Def. B.2),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. B.3),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. B.5), and
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. B.6).
@@
where the 1st step follows from Lemma 4.1, and all other steps follow from Fact A.9.
@@
Lemma C.3 (Gradient Reformulation, $\frac{\mathrm{d}\mathsf{Loss}(x)_{j_{0}}}{\mathrm{d}x}$).
@@
  • The column function $u(x)_{j_{0}}\in\mathbb{R}^{n}$ (Definitions B.1),
  • $\alpha(x)_{j_{0}}$ is a real number (Def. B.2),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. B.3),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. B.5),
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. B.6), and
@@
$\frac{\mathrm{d}\mathsf{Loss}(x)_{j_{0}}}{\mathrm{d}x}=$ $~\widetilde{\mathsf{A}}_{j_{0}}^{\top}(s(x)_{j_{0}}\circ A_{3}\beta(x)_{j_{0}})$
- $~\widetilde{\mathsf{A}}_{j_{0}}^{\top}s(x)_{j_{0}}\langle s(x)_{j_{0}},A_{3}\beta(x)_{j_{0}}\rangle.$
@@
$~\frac{\mathrm{d}\mathsf{Loss}(x)_{j_{0}}}{\mathrm{d}x}$
$=$ $~\sum_{i_{0}\in[d]}\frac{\mathrm{d}\mathsf{Loss}(x)_{j_{0},i_{0}}}{\mathrm{d}x}$
@@
Lemma C.4 (Gradient Reformulation, $\frac{\mathrm{d}\mathsf{Loss}(x)}{\mathrm{d}x}$, Formal version of Lemma 4.1).
@@
  • The column function $u(x)_{j_{0}}\in\mathbb{R}^{n}$ (Definitions B.1),
  • $\alpha(x)_{j_{0}}$ is a real number (Def. B.2),
  • $s(x)_{j_{0}}$ is an arbitrary element in $\mathbb{R}^{n}$ (Def. B.3),
  • $\ell(x)_{j_{0},i_{0}}$ is a real number (Def. B.5),
  • $\mathsf{Loss}(x)_{j_{0},i_{0}}$ is a real number (Def. B.6),
@@
C.2 Time Complexity for Computing $s(x)$ and $v(y)$ Functions
In this section, we use the vector and matrix multiplication time complexity from Definition A.2 and Fact A.3 to analyze the complexity of computing $s(x)$ and $v(y)$.
Lemma C.5.
Pick $s(x)$ and $v(y)$ from Def. B.3 and Def. B.4, then it costs $O(\mathcal{T}_{\mathrm{mat}}(n,d,d)+\mathcal{T}_{\mathrm{mat}}(n,d,n))$ time to get $s(x)$, and it costs $\mathcal{T}_{\mathrm{mat}}(n,d,d)$ time to get $v(y)$.
@@ C.2 Time Complexity for Computing $s(x)$ and $v(y)$ Functions
C.3 Time Complexity for Computing $\ell(x)$ Functions
In this section, we use the vector and matrix multiplication time complexity from Definition A.2 and Fact A.3 to analyze the complexity of computing $\ell(x)$.
Lemma C.6.
We have $\ell(x)$ from Def. B.5, then it costs $\mathcal{T}_{\mathrm{mat}}(n,n,d)+O(nd)$ to calculate $\ell(x)$.
@@ C.3 Time Complexity for Computing $\ell(x)$ Functions
Then, it requires $O(nd)$ time to get s(x)v(y)Es(x)v(y)-E.
@@
C.4 Time Complexity for Computing $\beta(x)$ and $\gamma(x)$ Functions
In this section, we use the vector and matrix multiplication time complexity from Definition A.2 and Fact A.3 to analyze the complexity of computing $\beta(x)$ and $\gamma(x)$.
Lemma C.7.
@@ C.4 Time Complexity for Computing $\beta(x)$ and $\gamma(x)$ Functions
C.5 Time Complexity for Computing the Gradient of RoPE Attention
Theorem C.8 (RoPE attention gradient computation time complexity, Restatement of Theorem 4.2).
@@ C.5 Time Complexity for Computing the Gradient of RoPE Attention
    We need time $O(\mathcal{T}_{\mathrm{mat}}(n,d,d)+\mathcal{T}_{\mathrm{mat}}(n,d,n))$ for $s(x),v(y)$ from Lemma C.5.
@@
    We need time $O(\mathcal{T}_{\mathrm{mat}}(n,n,d)+\mathcal{T}_{\mathrm{mat}}(n,d,d))$ for $\ell(x)$ from Lemma C.6.
@@
    We need time $O(\mathcal{T}_{\mathrm{mat}}(n,n,d))$ for $\beta(x)$ from Lemma C.7.
@@
    We need time $O(n^{2})$ for $\gamma(x)$ from Lemma C.7.
@@
Appendix D Low Rank Approximation of RoPE Attention
This section presents the fast running time using the low-rank approximations where the low-rank matrices are generated from the polynomial method (see Lemma A.1).
D.1 Approximate $s$ Using Low Rank Approximation
@@
Lemma D.1 (Low Rank Approximate $s(x)$).
@@
D.2 Approximate $\ell$ Using Low Rank Approximation
@@
Lemma D.2 (Low Rank Approximate $\ell(x)$).
Let $d$ equal $O(\log n)$. Suppose we can use $O(\log n)$ bits to write every entry in $E,v(y)\in\mathbb{R}^{n\times d}$. Define the $\ell(x)\in\mathbb{R}^{n\times d}$ as specified in Def. B.5. Then, we have $U_{1},V_{1}\in\mathbb{R}^{n\times k_{1}}$ such that $\|U_{1}V_{1}^{\top}v(y)-E-\ell(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n).$
@@
where the 1st is because of Def. B.5, 2nd step is based on the distributive law, and 3rd step is due to Lemma D.1. ∎
D.3 Approximate $\beta$ Using Low Rank Approximation
@@
Lemma D.3 (Low Rank Approximate $\beta(x)$).
Let $k_{2}=n^{o(1)}$. We define $\ell(x)\in\mathbb{R}^{n\times d}$ based on Def. B.5, and $v(y)\in\mathbb{R}^{n\times d}$ based on Def. B.4. We suppose $\beta(x)$ is equal to $v(y)\ell(x)^{\top}$, which is an $n\times n$ matrix. Let $U_{2},V_{2}\in\mathbb{R}^{n\times k_{2}}$ such that $\|U_{2}V_{2}^{\top}-\beta(x)\|_{\infty}\leq\epsilon/\operatorname{poly}(n)$. In $n^{1+o(1)}$ time, we can get $U_{2},V_{2}$.
@@
By Lemma D.2, $U_{1}V_{1}^{\top}v(y)-E$ approximately equals to $\ell(x)$.
@@
where the first step is based on the definition of $\beta(x)$ and $\widetilde{\beta}(x)$, the second step is due to the distributive law, and the third step derives from Lemma D.2.
@@
D.4 Approximate $\gamma$ Using Low Rank Approximation
@@
Lemma D.4 (Low Rank Approximate $\gamma_{1}(x)$).
@@
Then, we have the following by Fact A.5.
@@
where the first inequality is because of the def. of $\gamma_{1}(x)$, the second equality is due to the def. of $U_{3},V_{3}$, the third equality is based on Fact A.5, the fourth equality is due to the def. of $\widetilde{s}(x)$ and $\widetilde{\beta}(x)$, the fifth equality is due to simple arithmetic, the sixth inequality is because of the triangle inequality, and the seventh inequality derives from Lemma D.1 and Lemma D.3. ∎
@@
Lemma D.5 (Low Rank Approximate $\gamma_{2}(x)$).
@@
Then, we define $\widetilde{\rho}(x)_{j_{0}}$ as the inner product of $\widetilde{s}(x)_{j_{0}}$ and $\widetilde{\beta}(x)_{j_{0}}$, and by Fact A.9, we have $\widetilde{\rho}(x)_{j_{0}}=(U_{1}V_{1})_{j_{0},*}\cdot(U_{2}V_{2})_{j_{0},*}^{\top}$
@@
D.5 Fast Computation in Almost Linear Time